In a time study of a manufacturing operation, 15 actual times were recorded for the duration of producing a product for a skilled worker who has an experience of 10 years and for an unskilled worker who started to work one months ago: For the skilled worker, the average time observed to complete a product was 12 minutes. The performance rating applied to the observed worker was 1.10 and the allowance during a 10-hour shift was 30% or 180 minutes. For the unskilled worker, the average time observed to complete a product was 15 minutes. The performance rating applied to the observed worker was 0.85 and the allowance during a 10-hour shift was 10% or 60 minutes. Compute the labor standard times for skilled workers and unskilled workers. Discuss your results by comparing labor standards of skilled and unskilled workers.
Answer #1

Skilled worker

Average observed time = 12 minutes

Performance rating = 1.10

Allowance factor = 30%

Normal time = Average observed time * performance rating

= 12*1.10

= 13.2 minutes

Standard time = Normal time / (1 - Allowance factor)

= 13.2 / (1 - 30%)

= 18.86 minutes

--------------------

Unskilled worker

Average observed time = 15 minutes

Performance rating = 0.85

Allowance factor = 10%

Normal time = Average observed time * performance rating

= 15*0.85

= 12.75 minutes

Standard time = Normal time / (1 - Allowance factor)

= 12.75 / (1 - 10%)

= 14.17 minutes

--------------------------------------------

Comparing the labor standards of skills and unskilled workers, we see that the standard time of unskilled worker is much lower than that of the skilled worker.